OpenStudy (anonymous):

The functions f and g are defined by f:x ---> 3x+2 x all Real numbers g:x ---> {6 over 2x+3} " xnotequalto -1.5

OpenStudy (anonymous):

find the value of x for which fg(x) = 3

OpenStudy (anonymous):

does x=3?

OpenStudy (anonymous):

f(g(x)) = 3(6/(2x+3)) +2 = 18/(2x+3) + 2 therefore 18/(2x+3) + 2 = 3 so 18/ (2x+3) = 1. so 2x+3 = 18 so 2x = 15 so x = 7.5

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

(ii) says sketch, in a single diagram, the graphs of y=f(x) and y=f^-1(x), making clear the relationship between the two graphs

OpenStudy (anonymous):

is f^-1 y=x-2 over 3?

OpenStudy (anonymous):